Fermi National Laboratory
 

Accelerator Update for June 20, 2007 - June 22, 2007

 

Wednesday June 20

The day shift began with Operations monitoring the Recycler stash (288.3E10), with the Antiproton Source (Pbar) stacking (18.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations put the Tevatron (TeV) into shot setup at 8:16 AM.

At 9:32 AM, a Linac Klystron RF station (KRF5) tripped off. After eight minutes it reset.

At 9:55 AM, KRF5 tripped off. After five minutes it reset.

Operations established store 5505 at 10:24 AM with an initial luminosity of 201.07E30.

Operations established beam to Switchyard (SY) and all the way down to the Meson MTest area at 1:07 PM. MTest won’t have an experiment in it until next week.

At 2:33 PM, the NuMI beam permit tripped off due to a bad beam position. It reset.

A Booster expert began tuning the machine at 3:12 PM.

At 3:58 PM, Operations reported that there was a problem with the Recycler RF. Experts investigated.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons twice during the day shift.

At 5:38 PM, a SY expert put MTest into its low energy pion mode at 16 GeV. When he tried to switch to the 2 GeV energy, a problem with a power supply (F:MT4W) halted all further testing at 6:17 PM. Operations disabled the MTest beam at 6:26 PM and left a message for the power supply experts to check it out on Thursday.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Thursday June 21

The midnight shift began with Operations monitoring store 5505 (57.54E30) and the stash (217.5E10), with Pbar stacking (24.9E10), and with NuMI and MiniBooNE taking beam.

At 4:31 AM, the NuMI beam permit tripped off due to a MI RF problem. Everything reset.

At 7:08 AM, a Booster vacuum burst caused BRF7 and 8 to trip as well as some Booster vacuum pumps. Everything reset.

Operations reported at 7:58 AM that the Recycler had stashed three times during the midnight shift.

Operations terminated store 5505 at 9:01 AM.

Operations prepared CDF for access at 9:14 AM.

Operations established beam to MTest at 9:20 AM for expert tuning and beam studies.

CDF came out of access at 11:20 AM.

The Neutron Therapy Facility reported at 11:49 AM that they had finished treating their last patient for the day.

Operations put the TeV through a dry squeeze at 12:43 PM.

Operations put the TeV into shot setup at 1:22 PM.

Operations established store 5506 at 3:07 PM with an initial luminosity of 220.07E30.

Experts replaced at 3:42 PM the beam monitor card at MI-30 that was so sensitive to the humidity.

MTest experts completed their studies at 3:43 PM. Next Wednesday, 6/27/07, a new experiment will begin their run.

Operations reported at 3:55 PM that the Recycler had stashed twice during the day shift.

A Recycler expert reported at 5:42 PM that he had completed the new Recycler vertical damper installation.

At 10:04 PM, the Booster High Level RF monitor tripped all of its interlock bits. Operators investigated and found a bad power supply for a NIM crate. Operators changed the power supply. Beam resumed at 11 PM.

At 11:37 PM, a MI kicker (I:KPS1A) tripped off. It reset.

Operations reported at 11:58 PM that the Recycler had stashed three times during the evening shift.

Friday June 22

The midnight shift began with Operations monitoring store 5506 (89.97E30) and the stash (166.4E10), with Pbar stacking (24.7E10), and with NuMI and MiniBooNE taking beam.

At 12:18 AM, a Recycler RF amplifier (AR5KW2) tripped. Operations contacted an expert and then at 1:58 AM replaced the RF amp.

At 2:13 AM, the NuMI horn tripped off. Operators reset it twice, but both times it tripped on a ground fault indication. An expert said this was a flaking paint problem. Operators followed the procedure to burn off the flake. NuMI beam resumed at 2:21 AM.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow 15 minutes to Pbar and Recycler experts for a study and tuning today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/22/2007    email Fermilab